# Topological Devil’s staircase in a constrained kagome Ising antiferromagnet

### v1

@@ Abstract
In frustrated models, the impossibility of satisfying every interaction may lead to numerical instabilities in CTMRG at low temperatures and prevent convergence [45, 63]. A helpful strategy is to choose a different local tensor, defining a TN corresponding to the same partition function, but such that the constraints are enforced locally within the tensor. To construct this local tensor, we follow the approach of Ref. [45] where the Boltzmann weight of each spin configuration is factorized over overlapping clusters. The TN is then built from these local Boltzmann weights by imposing the matching of overlapping spins in adjacent clusters. Here, we chose kagome stars as clusters. Each pair of corner-sharing triangles is shared between two clusters. This would normally give rise to a triangular TN, but, thanks to the redundancy of the matching conditions, it is enough to impose them in two out of three directions, giving rise to the square TN illustrated in Fig. EM2b. This results in
$a_{i_{1},i_{2},i_{3},i_{4}}=\sum_{k=1}^{2^{11}}\delta_{i_{1},i_{1}(k)}\delta_{i_{2},i_{2}(k)}\delta_{i_{3},i_{3}(k)}\delta_{i_{4},i_{4}(k)}e^{-E_{k}/T},$ (EM1)
@@
where $E_{k},k\in\{1,\dots,2^{11}\},$ are the energies of the cluster configurations, and where functions $i_{a}(k)$ encode the configurations of the two overlapping triangles in each direction. Note that we take advantage of the 2-to-1 mapping between spin and dual-bond configurations to divide the bond dimension by two: thanks to the global spin-flip symmetry of the model, it is sufficient to keep track of only the dual-bond configurations. Since there are four possible dual bond configurations for each triangle (Fig. EM2c), the TN has bond dimension 16.
@@
As mentioned in the main text, the ground states of the strings phase are dominated by a partially ordered family with antiferromagnetic long-range order on the dense rows and quasi-long-range order on the sparse rows. Thus, the strings phase breaks the $\mathbb{Z}_{3}$ rotational symmetry and the $\mathbb{Z}_{2}$ spin-flip symmetry at low temperatures. We numerically detect the restoration of these symmetries as follows. First, the initial tensors for the environment are chosen such that the $\mathbb{Z}_{3}$ symmetry is slightly broken by favoring a certain orientation of the crosses (this is the equivalent of slightly favoring up spins in the initial condition for a TN simulation of the 2D square lattice Ising model: it only selects a finite magnetization if there is spontaneous symmetry breaking.). This allows us to compute the $\mathbb{Z}_{3}$ order parameter using a local tensor which detects whether a particular dual bond is favored:
$\Psi_{\mathbb{Z}_{3}}$ =|$\sigma_{1}(\mathbf{0})\sigma_{2}(\mathbf{0})+\sigma_{2}(\mathbf{0})\sigma_{3}(\mathbf{0})e^{2i\pi/3}\right.\right.$ (EM2)
$+\left.\left.{\sigma_{3}(\mathbf{0})\sigma_{1}(\mathbf{0})e^{-2i\pi/3}}\right>_{T}\right|.$
Here, $\sigma_{a}(\mathbf{r}_{n_{1},n_{2}})$ is the spin at position $\mathbf{R}_{n_{1},n_{2},a}=\mathbf{r}_{n_{1},n_{2}}+\boldsymbol{\tau}_{a}$, with $\mathbf{r}_{n_{1},n_{2}}=n_{1}\mathbf{a}_{1}+n_{2}\mathbf{a}_{2}$, where $\mathbf{a}_{1}=(1,0)$ and $\mathbf{a}_{2}=\left(-1/2,-\sqrt{3}/2\right)$, and the vectors $\boldsymbol{\tau}_{1}=0$, $\boldsymbol{\tau}_{2}=\mathbf{a}_{1}/2$, and $\boldsymbol{\tau}_{3}=(\mathbf{a}_{1}+\mathbf{a}_{2})/2$ specify the positions of the three sites within each unit cell of the kagome lattice.
@@
$\Psi_{\mathbb{Z}_{2}}$ $=\lim_{r\rightarrow\infty}\sqrt{|\left<\sigma_{1}(\mathbf{0})\sigma_{1}(r\mathbf{e}_{x})\right>_{T}|}.$ (EM3)
Figure EM3: CTMRG results ($\chi=100$) for the temperature dependence of the order parameters for the $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$ broken symmetries, with infinite $J_{1}$ and $J_{3}$.
The results for the order parameters as a function of temperature are shown in Fig. EM3. The $\mathbb{Z}_{2}$ order parameter drops to zero at $T_{c}^{(1)}=1.78J_{2}$, while the $\mathbb{Z}_{3}$ order parameter remains non-zero up to the final transition temperature $T_{c}^{(\infty)}=3.99J_{2}$: the intermediate phases are nematic. As the temperature increases, $\Psi_{\mathbb{Z}_{3}}$ decreases through a sequence of discontinuous jumps until it vanishes when the density of C lines reaches its high-temperature value of $2/3$. Indeed, the $\mathbb{Z}_{3}$ order parameter is related to $n_{C}$ because of the constrained imposed on nearest-neighbor triangles by the limit of infinite $J_{1}$, together with the reflection symmetry
$\sigma_{1}(\mathbf{r})\sigma_{3}(\mathbf{r})\right>_{T}=\left<\sigma_{2}(\mathbf{r})\sigma_{3}(\mathbf{r})\right>_{T},$ (EM4)
as follows:
$\Psi_{\mathbb{Z}_{3}}=1/2-3\left<\sigma_{1}(\mathbf{0})\sigma_{2}(\mathbf{0})\right>_{T}/2=1/2-3n_{C}/4.$ (EM5)
@@
An often overlooked aspect of CTMRG is its ability to efficiently sample large snapshots of systems according to the equilibrium Gibbs distribution, as demonstrated by Ueda et al. in 2005 [46] for the case of the square-lattice Ising model. The key idea is to sequentially sample spins in a rectangular region. The first spin is sampled at random from the distribution given by the environment. The conditional probability distribution for the next spin given the already sampled ones can be computed as a TN contraction whose cost grows linearly with the snapshot size.
@@
S2 Dependence of results on bond dimension
@@
We suspect the $J_{1},J_{3}\rightarrow\infty$ constrained limit of the kagome Ising antiferromagnet to be critical at all temperatures. In the low- and high temperature limits, we have strong evidence. If $T\rightarrow 0$, spins on the disordered sub-lattice are correlated as in the nearest neighbor triangular-lattice Ising antiferromagnet, which is critical. In the $T\rightarrow\infty$ limit of the constrained model, one obtains the $J_{1}-J_{3}$ kagome Ising antiferromagnet, whose ground-state correlations decay algebraically with a critical exponent $\eta=0.21$ (data to appear in the follow-up work mentioned in the conclusion). We expect the system to remain critical at all intermediate temperatures. Additional evidence for this is given by the unbounded growth, at all temperatures, of the correlation length, as computed from the eigenvalues of the row-to-row transfer matrix (data not shown).
As a consequence of criticality, observables computed with CTMRG on the constrained model display strong bond dimension dependence. If the ratio of densities $n_{C}/n_{A}$ is plotted with different values of $\chi$ (Fig. S2a), we can observe that the plateau with $n_{C}/n_{A}=4$ is present only if $\chi\geq 70$, while plateaus with smaller integer ratios are present even at lower bond dimensions. To understand why, we recall that spin correlations in the quantized phases are modulated by a characteristic length equal to the average distance between A strings: $\bar{l}_{1}=1/n_{A}\approx n_{C}/n_{A}+3$. This distance increases with the value of the integer ratio, which means that the higher-temperature plateaus can only be accurately described by CTMRG at larger bond dimensions. Note that this numerical correlation length is much larger than $\overline{l}_{1}$.
@@
Finally, we also discuss the presence of points on the boundary between the phases, for example at $T/J_{2}=3.64$ which seem to violate the quantization of $n_{C}/n_{A}$ to integer values. A finer grained calculation around this temperature (Fig. S2) reveals that these points belong to the crossover between adjacent plateau phases. The width of the crossover is narrowed as $\chi$ is increased, exactly as in a finite system near a first-order phase transition. The numerical results support the hypothesis that in the thermodynamic limit ($\chi\rightarrow\infty$) $n_{C}/n_{A}$ is exactly quantized to integer values.
S3 Approximate calculations of the magnetic structure factor
@@
[…]
@@
$\sum_{(a,b)\in\{1,2\}^{2}}e^{i\mathbf{q}\cdot(\boldsymbol{\tau}_{a}-\boldsymbol{\tau}_{b})}G_{a,b}(\mathbf{q})$ $=4\pi\left(\sum_{m_{2}\in\mathbb{Z}}\delta\left(q_{y}-\frac{4\pi m}{\sqrt{3}}\right)\right)\left(\sum_{k=1}^{\infty}\frac{4}{\pi^{2}(2k-1)^{2}}\left[\sum_{u\in\mathbb{Z}}\exp\left\{\frac{i}{2}\left(q_{x}+2\pi+\frac{\pi(2k-1)}{\bar{l}_{1}}\right)u\right\}\right.\right.$ (S12)
@@
If $\bar{l}_{1}$ is incommensurate with the underlying lattice spacing, the set of peak positions $\mathbf{Q}^{p=1}_{(m_{1},m_{2},n)}$ is in principle dense in the horizontal lines defined by $q_{y}=\frac{4\pi}{\sqrt{3}}m_{2}$, a common situation for one-dimensional incommensurate systems [49, 68]. However, the prefactor $1/(2k-1)^{2}$ strongly suppresses peaks with large values of $k$, and peaks in the vicinity of a given peak at a small value of $k$ would correspond to very large values of $k$ shifted back by a large value of $m_{1}$, and in practice they are invisible. The one-dimensional model of uniformly spaced domain walls thus accurately predicts that peaks will be visible at $\frac{\pi(2k-1)}{\bar{l}_{1}}$ for small $k$. However, it does not explain why only the three peaks with $0\leq k\leq 2$ can be observed. In part, the vanishing intensity of peaks with large $|k|$ can be attributed to the randomness of the domain wall positions, also responsible for smoothening the Bragg peaks. This hypothesis was confirmed numerically by calculating the MSF of a one-dimensional model where the distance between adjacent domain walls is a random variable of average $\bar{l}_{1}$ and variance $(\delta\bar{l}_{1})^{2}$ (data not shown). However, the intensity asymmetry which leads to the peak with $k=-1$ but not the one with $k=2$ being extinguished is not simply a result of fluctuations. As discussed in more detail later, this results from the correlation with spins on sparse rows.
@@
$\mathbf{Q}^{p}_{(m_{1},m_{2},k)}=\begin{cases}\left(4\pi m_{1}+2\pi+\frac{\pi(2k-1)}{\bar{l}_{1}},\frac{4\pi m_{2}}{\sqrt{3}}\right)\text{, if p is odd}\\ \left(4\pi m_{1}+2\pi+\frac{\pi(2k)}{\bar{l}_{1}},\frac{4\pi m_{2}}{\sqrt{3}}\right)\text{, if p is even}\end{cases}.$ (S15)
@@
S4 Additional ground states of the strings phase

### v2

@@ Abstract
In frustrated models, the impossibility of satisfying every interaction may lead to numerical instabilities in CTMRG at low temperatures and prevent convergence [47, 69]. A solution is to choose a different TN formulation, such that the constraints are enforced within a single tensor. To construct this local tensor, we follow the approach of Ref. [47] where the Boltzmann weight is factorized over overlapping clusters, with the contraction of the indices in the network imposing the matching conditions. Here, our clusters are kagome stars. Each pair of corner-sharing triangles is shared between two clusters. This would normally give rise to a triangular TN, but, thanks to the redundancy of the matching conditions, two out of three directions suffice, giving rise to the square TN illustrated in Fig. EM2b.
@@
We further take advantage of symmetry to divide the bond dimension by two: since configurations related by a global spin-flip have the same energy, it is sufficient to keep track of the position of the ferromagnetic bonds (dual bond configurations). Since there are four possible dual bond configurations per triangle (Fig. EM2c), the TN has bond dimension 16.
@@
As mentioned in the main text, the ground states of the strings phase are dominated by a partially ordered family with antiferromagnetic long-range order on the dense rows and quasi-long-range order on the sparse rows. Thus, the strings phase breaks the $\mathbb{Z}_{3}$ rotational symmetry and the $\mathbb{Z}_{2}$ spin-flip symmetry at low temperatures. We numerically detect the restoration of these symmetries as follows. First, the initial tensors for the environment are chosen such that the $\mathbb{Z}_{3}$ symmetry is slightly broken by favoring a certain orientation of the crosses (this is the equivalent of slightly favoring up spins in the initial condition for a TN simulation of the 2D square lattice Ising model: it only selects a finite magnetization if there is spontaneous symmetry breaking.).
This allows us to compute the $\mathbb{Z}_{3}$ order parameter using a local tensor which detects whether a particular dual bond is favored:
[…]
$+\left.\left.{\sigma_{3}(\mathbf{0})\sigma_{1}(\mathbf{0})e^{-2i\pi/3}}\right>_{T}\right|,$
where $\{\sigma_{1}(\mathbf{R}),\sigma_{2}(\mathbf{R}),\sigma_{3}(\mathbf{R})\}$ are the three spins at the vertices of a triangle centered around $\mathbf{R}$. This $\mathbb{Z}_{3}$ order parameter is related to the density of of C strings by $\Psi_{\mathbb{Z}_{3}}=1/2-3n_{C}/4$, which results from reflection symmetry and the constraints imposed by infinite $J_{1}$.
@@
$\Psi_{\mathbb{Z}_{2}}$ $=\lim_{r\rightarrow\infty}\sqrt{|\left<\sigma_{1}(\mathbf{0})\sigma_{1}(r\mathbf{e}_{x})\right>_{T}|}.$ (EM2)
Figure EM3: CTMRG results ($\chi=100$) for the temperature dependence of the order parameters for the $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$ broken symmetries, with infinite $J_{1}$ and $J_{3}$. Only temperatures where the results are converged in bond dimension are shown.
The numerical calculation of both order parameters are shown in Fig. EM3. $\mathbb{Z}_{2}$ is restored at the first phase transition ($T_{c}^{(1)}=1.78J_{2}$), while the $\mathbb{Z}_{3}$ symmetry is only restored at the accumulation point ($T_{c}^{(\infty)}=3.99J_{2}$), which means that the phases between these temperatures are nematic.
@@
An often overlooked aspect of CTMRG is its ability to efficiently sample large snapshots of systems according to the equilibrium Gibbs distribution, as demonstrated by Ueda et al. in 2005 [48] for the case of the square-lattice Ising model. The key idea is to sequentially sample spins in a rectangular region. The first spin is sampled at random from the distribution given by the environment. The conditional probability distribution for the next spin given the already sampled ones can be computed as a TN contraction whose cost grows linearly with the snapshot size. The sampling of system snapshots using CTMRG may also have applications in hybrid Tensor-Network Monte-Carlo schemes [73, 74].
@@
S2 Real-space correlations
As briefly mentioned in the main text, the ability to sample system snapshots permits the calculation of correlations between any two spins. In Fig. S2, the correlations between spins in different rows and columns of the Kagome lattice are shown at different temperatures. The highly anisotropic nature of spin-spin correlations is evident in the fact that vertically separated correlations are positive and monotonously decreasing, while horizontal correlations are modulated by commensurate or incommensurate wave-vectors (depending on temperature). Looking at the plots in a log-log scale, correlations can be noticed to decay algebraically with distance at all temperatures.
The algebraic correlations were further characterized by the $\eta$ critical exponent, defined by the asymptotic expression
$G(r)\sim f_{\text{osc.}}(r)r^{-\eta}\Rightarrow\log G(r)\sim\log\left\{f_{\text{osc.}}(r)\right\}-\eta\log r,$ (S3)
where $1\leq f_{\text{osc.}}(r)\leq 1$ is a an oscillating function, zero on average, and which does not decay with distance. In the non-oscillating case ($f_{\text{osc.}}(r)=1$) $\eta$ is determined through a linear fit to the log-log correlation curve. In the oscillating case, due to the challenges resulting from incommensurability, we estimate $\eta$ from the slope of the upper convex hull of the absolute value of correlations in log-log scale: at large distances, the upper convex hull follows the amplitude of oscillations, equal to evaluating $G(r)$ in a sequence of points where $|f_{\text{osc.}}(r)|\approx 1$. The determination of the $\eta$ exponent reveals that correlations between spins on dense rows decay faster at higher temperatures, while the correlations between spins on sparse rows decay faster at lower temperatures.
Figure S2: Real-space correlations between spins on different curves of the Kagome lattice at four temperatures, corresponding to phases with different values of $n_{C}/n_{A}$. Each color in the correlation plots corresponds to a row or column of the Kagome lattice, as prescribed by the illustration on the top-left. 1st row: Correlation function in the vertical direction. 2nd row: Correlation function in the vertical direction plotted in a log-log scale. The critical exponent $\eta$ is determined from a linear fit. 3rd row: Correlation function in the horizontal direction. On the leftmost panel, the exact spin-spin correlations of the Triangular Ising antiferromagnet plotted in black [76]. 4th row: Correlation functions plotted in log-log scale. Due to incommensurability, the critical exponent $\eta$ is determined from the slope of the convex hull of the log-log plots. Correlations were calculated by averaging over more than 2700 snapshots of size $40\times 40$. Error bars indicate statistical uncertainties (whenever error bars are not visible, they are smaller than the data-point markers).
S3 Dependence of results on bond dimension
@@
As discussed in section S2, our numerical results show that the constrained limit of the Kagome Ising antiferromagnet is critical at all temperatures. As a consequence of criticality, observables computed with CTMRG on the constrained model display strong bond dimension dependence. If the ratio of densities $n_{C}/n_{A}$ is plotted with different values of $\chi$ (Fig. S3a), we can observe that the plateau with $n_{C}/n_{A}=4$ is present only if $\chi\geq 70$, while plateaus with smaller integer ratios are present even at lower bond dimensions. To understand why, we recall that spin correlations in the quantized phases are modulated by a characteristic length equal to the average distance between A strings: $\bar{l}_{1}=1/n_{A}\approx n_{C}/n_{A}+3$. This distance increases with the value of the integer ratio, which means that the higher-temperature plateaus can only be accurately described by CTMRG at larger bond dimensions. Note that the numerical correlation length is much larger than $\overline{l}_{1}$.
@@
Finally, we also discuss the presence of points on the boundary between the phases, for example at $T/J_{2}=3.64$, which seem to violate the quantization of $n_{C}/n_{A}$ to integer values. A finer grained calculation around this temperature (Fig. S3) reveals that these points belong to the crossover between adjacent plateau phases. The width of the crossover is narrowed as $\chi$ is increased, exactly as in a finite system near a first-order phase transition. The numerical results support the hypothesis that in the thermodynamic limit ($\chi\rightarrow\infty$) $n_{C}/n_{A}$ is exactly quantized to integer values.
S4 Approximate calculations of the magnetic structure factor
@@
Figure S4: Plots of the MSF of the $J_{1},J_{3}\rightarrow\infty$ constrained model along the $(h,0,0)$ direction.
@@
[…]
@@
If $\bar{l}_{1}$ is incommensurate with the underlying lattice spacing, the set of peak positions $\mathbf{Q}^{p=1}_{(m_{1},m_{2},n)}$ is in principle dense in the horizontal lines defined by $q_{y}=\frac{4\pi}{\sqrt{3}}m_{2}$, a common situation for one-dimensional incommensurate systems [51, 80]. However, the prefactor $1/(2k-1)^{2}$ strongly suppresses peaks with large values of $k$, and peaks in the vicinity of a given peak at a small value of $k$ would correspond to very large values of $k$ shifted back by a large value of $m_{1}$, and in practice they are invisible. The one-dimensional model of uniformly spaced domain walls thus accurately predicts that peaks will be visible at $\frac{\pi(2k-1)}{\bar{l}_{1}}$ for small $k$. However, it does not explain why only the three peaks with $0\leq k\leq 2$ can be observed. In part, the vanishing intensity of peaks with large $|k|$ can be attributed to the randomness of the domain wall positions, also responsible for smoothening the Bragg peaks. This hypothesis was confirmed numerically by calculating the MSF of a one-dimensional model where the distance between adjacent domain walls is a random variable of average $\bar{l}_{1}$ and variance $(\delta\bar{l}_{1})^{2}$ (data not shown). However, the intensity asymmetry which leads to the peak with k=1k=-1 but not the one with $k=2$ being extinguished is not simply a result of fluctuations. As discussed in more detail later, this results from the correlation with spins on sparse rows.
@@
[…]
@@
S5 Additional ground states of the strings phase
@@
S6 System-spanning strings from the $J_{1},J_{3}\rightarrow\infty$ constraints
We have mentioned in the main text that the allowed configurations of the $J_{1},J_{3}\rightarrow\infty$ limit of the Kagome Ising antiferromagnet are associated to the trajectories of three kinds of system-spanning strings:
  • A strings, drawn by connecting the centers of crosses arrows where arrow directions converge.
  • B strings between pairs of A strings, drawn by connecting the centers of hexagons where arrow directions diverge.
  • C strings, consisting of Double Domain Walls (DDW’s) with internal excitations (changes in arrow directions) and decorating B lines.
In this section, we first illustrate how the $J_{1},J_{3}\rightarrow\infty$ limit leads to local constraints, and then how those constraints lead to A and C strings which must be system-spanning.
Let us start from the $J_{1}\rightarrow\infty$ limit. Because $J_{2}$ and $J_{3}$ are not competing with $J_{1}$, this limit imposes a simple 2-up 1-down, 2-down 1-up (UUD-DDU) rule on every nearest-neighbor triangle of the kagome lattice. As illustrated in the End Matter (Fig. EM2), this constraint is equivalent to having exactly one dimer touching the center of every triangle. The limit $J_{3}\rightarrow\infty$ plays a similar role, imposing the UUD-DDU rule on three independent triangular sublattices of the model. On the level of a kagome star, these two constraints conspire to give rise to all the configurations illustrated in Fig. EM1. These configurations can be understood as local tiles or dominoes, which overlap and impose constraints, and must be matched to realize configurations of the kagome lattice model [47, 41]. We now show how these constraints give rise to system spanning strings.
We begin with the case of the A strings. Our starting tile is a single crossed hexagon, and we seek to determine what are the allowed values for the surrounding spins. Since the energy of every $J_{1}$ and $J_{3}$ triangle is minimized, whenever two spins which interact with $J_{1}$ or $J_{3}$ are ferromagnetically aligned the value of the third spin on the triangle is set by constraints. Repeatedly applying this principle as shown in FIG. S8, we conclude that beneath (and above, by inversion symmetry) a crossed hexagon there must be another crossed hexagon. Successive application of this reasoning leads to the conclusion that crossed hexagons are always a part of system-spanning strings (see also the proof in Appendix of [41]). A further consequence of this is that in the constrained system all crosses must have the same orientation: crosses with different orientations will propagate on average in different directions and must eventually intersect, which is not allowed.
Figure S8: Starting from a hexagon with a cross, the constraints imposed by the $J_{1},J_{3}\rightarrow\infty$ force the existence of another cross in one of the two adjacent hexagons beneath (by symmetry also above) it. At each step of the illustration the value of a spin is determined by the existence of two identical spins in a $J_{1}$ or $J_{3}$ triangle (highlighted in gray). The value of one of the spins may be freely chosen, leading to two branches for the diagram. This process may be continued indefinitely, leading to a system-spanning A strings which must not intersect.
We now move onto the case of the C strings. We prove that if there is at least one A string in the system, and that it has the orientation shown in FIG. S8, then C strings must be oriented and non-crossing. We first prove some statements about C-strings and dimers, before restarting from the $J_{3}\rightarrow\infty$ constraint to conclude the proof.
The presence of one C string can be detected by the existence of a ferromagnetic horizontal bond (vertical dimer). Therefore, in order to prove that C strings cannot terminate it suffices to show that beneath (or above) a vertical dimer there must be another vertical dimer. We will need a useful intermediate result, which we now derive: adjacent vertical dimers must be between two crossed hexagons. The derivation is shown in FIG. S9 and results from the successive application of the minimization of $J_{1}$ and $J_{3}$ interactions in different triangles. Using the assumption about the orientation of crosses in A strings, this means that consecutive vertical dimers are not allowed.
Figure S9: In order to respect the $J_{1},J_{3}\rightarrow\infty$ constraints, two consecutive ferromagnetic bonds along a row must be between two diagonally oriented crosses.
Now, recall that the constraints imposed by infinite $J_{3}$ in one of the two triangular sub-lattices involving the dense rows force the spins in this sub-lattice to be ground states of the first-neighbor triangular Ising antiferromagnet. Recall as well that those configurations can be mapped onto the trajectories of system-spanning strings which separate stripe-ordered domains [59, 72, 21]. An example of an allowed configuration on this sub-lattice is represented in FIG. S10.
Since, as shown before, consecutive vertical dimers are not allowed, there is no freedom in the value of spins on a second $J_{3}$ sub-lattice, except when they overlap with the string trajectories. Spins along the length of a string are not fully fixed by constraints, but regardless of the choice that is taken there will always be one vertical dimer adjacent to the string, which shows that strings of vertical dimers must never terminate. The remaining freedom of the spins on the strings is due to the internal freedom of the C strings.
Figure S10: Illustration of how the $J_{1},J_{3}\rightarrow\infty$ constraints lead to system-spanning C strings. a) The allowed configurations of one of the $J_{3}$ sub-lattices may be described by trajectories of system-spanning strings (shown as dark lines), as in the first-neighbor triangular Ising antiferromagnet [60, 59]. b) In the presence of at least one horizontal cross (oriented as in Fig. S8), consecutive vertical dimers are not allowed. Therefore, the spins on the other $J_{3}$ sub-lattice of dense rows are also fully determined except on top of the strings. c) There are several possible choices for the spins along the black string (highlighted with a yellow star). Once the value of those spins has been chosen, the position of the vertical dimers is determined. It would also be possible to draw the black string using a different sub-lattice choice, which would have lead to the gray string. The strings drawn with different sublattice choices are always separated by half of a lattice spacing. d) Applying the $J_{1}\rightarrow\infty$ rules, the presence of the vertical dimers leads to system-spanning C strings. In some cases the trajectory of the C string is not fully determined from the position of the vertical dimers, due to the remaining freedom in the spins highlighted with a yellow star.